### v1

@@ 2 Background
LLMs can be broadly divided into two categories: traditional LLMs and reasoning LLMs. Traditional LLMs, such as early versions of GPT or similar models, are primarily optimized for fluent language generation and general-purpose tasks. Their strength lies in producing coherent and contextually appropriate text. However, their capabilities in logical reasoning and structured problem-solving are limited. Reasoning LLMs represent a newer generation of models that are explicitly designed to perform structured reasoning tasks more effectively. These models consider additional training objectives, architectural innovations, or fine-tuning procedures that enhance their ability to perform logical inference, complex decision-making, and consistent multi-step problem-solving. Reasoning LLMs aim not only at linguistic fluency but also at improved logical accuracy and reliability in formal contexts.
@@ 3 Methodology
    Controls randomness; higher values yield more diverse outputs [13]. Set to 00 to prioritize consistency over creativity [14].
@@ 3.1 Segmentation into Law Snippets
A key challenge in instructing the LLMs was determining the optimal length for law snippets: overly long snippets risked losing critical information during formalization, whereas overly short ones hindered atom reuse. To address this, we instructed the models to split enumerations containing more than two elements into separate law snippets while preserving shorter ones intact.
@@ 3.3 Evaluation
We evaluate the generated rules across six dimensions: completeness (Q1), syntactic (Q2) and semantic correctness (Q3), deontic modality accuracy (Q4), precondition appropriateness (Q5), and meaningfulness/reuse of atom names (Q6).
@@
Consider, for instance, the following formalization:
@@
This initial check is crucial to prevent the LLM from achieving a high score merely by formalizing only the simplest aspects of a problem. Unlike subsequent evaluations, however, Q1 is evaluated once per law snippet, not once per generated rule.
@@
Hence, the question would be answered with false, and no further checks performed. Again, note that this output stems from an early variation of the prompt. Such an error did not occur in later iterations.
@@
It is important to note that a single law snippet may lead to the generation of multiple rules. Furthermore, Q1 is assessed based on the law snippet as a whole, taking into account all rules derived from it, whereas Q2 through Q6 are evaluated separately for each individual rule.
@@
[…]
@@ 4.1 Prompt Development
The prompt employed in our experiments was derived through a series of iterative refinements. Beginning with an initial two-shot learning prompt, successive modifications were made to enhance the clarity of the instructions and the quality of the generated outputs. These iterations involved the inclusion of additional guidance and examples to better align the model’s behavior with the desired output format. The final version utilizes a three-shot learning approach. Outputs are expected to conform to the XML structure shown in Listing 2, where each rule is encoded as plain text within a <Rule> element, without further structuring. Listing 1 provides the complete and final prompt used in our evaluation.
@@ 8.5.1 A Supplier must take the following actions to
<Paragraph paragraphLabel="8.1.1.c">
<Rules>
<Rule ruleLabel="tcpc.8.1.1.c.1">
complaintHandlingProcess(X) => [O]
relevantStaffAwareComplaintHandlingProcess(X)
</Rule>
<Rule ruleLabel="tcpc.8.1.1.c.2">
complaintHandlingProcess(X) =>
[O] relevantStaffAbleToHandleComplaint(X)
</Rule>
</Rules>
</Paragraph>
Listing 2: Expected output format of the formalization
@@ 4.2.2 Providing Only Previously Formalized Atoms
In a second variant, previously extracted atom names were provided collectively in a single user message, rather than replicating the entire prior dialogue history. For each new law snippet, three messages were sent to the LLM: (1) a system prompt, (2) a user message listing previously formalized atom names (cf. Listing 3), and (3) a user message containing the new law snippet to be formalized.
@@
Listing 3: user message including previous atom names
@@ 4.3 Fine-Tuning
In the present study, fine-tuning experiments were conducted with GPT-4o. Although the proprietary nature of GPT-4o does not allow direct access to the model weights, OpenAI provides fine-tuning capabilities for non-reasoning models via its platform.
@@
Epochs $3$ $3$ $3$
Batch Size $1$ $4$ $4$
LR Multiplier $2$ $1.5$ $1$
@@
Fine-tuning resulted in an improved success score after a single epoch of training under Configurations 2 and 3, relative to the baseline performance of the non-fine-tuned GPT-4o. However, subsequent training epochs led to a decline in performance, indicative of overfitting. Furthermore, the fine-tuned models failed to achieve performance levels comparable to OpenAI’s most recent reasoning models, which are not amenable to fine-tuning.
@@ 4.4 Two-Stage Pipeline
In the first stage, atom names were extracted from the legal texts. To this end, the LLMs were instructed to identify atom names alongside brief textual descriptions (Listing 4). Three illustrative examples were provided within the prompt, thus applying a three-shot learning strategy.
@@ 8.1.1 A Supplier must take the following actions to
Listing 4: Prompt for atom extraction (2 examples omitted)
@@
In this section, we compare our experimental findings with the results presented by Dragoni et al. [8]. To ensure methodological consistency, we restrict the comparison to the lower branch evaluation reported in their study. For this purpose, we employ the standard metrics of Precision and Recall, defined as follows:
Precision=TPTP+FP=Matched ItemsGenerated ItemsRecall=TPTP+FN=Matched ItemsGold Standard $\begin{array}[]{rcccc}\text{Precision}&=&\displaystyle\frac{\text{TP}}{\text{TP}+\text{FP}}&=&\displaystyle\frac{\text{Matched Items}}{\text{Generated Items}}\\ \text{Recall}&=&\displaystyle\frac{\text{TP}}{\text{TP}+\text{FN}}&=&\displaystyle\frac{\text{Matched Items}}{\text{Gold Standard Items}}\end{array}$
Here, the term Items refers to either atoms or rules, depending on the context of the evaluation.
@@ 4.5.1 Evaluation of Term Identification
Experiment Model Generated Atoms Matched Atoms Precision Recall F1
@@ 4.5.3 Identification of Rule Counterparts
Experiment Model Generated Rules Matched Rules Precision Recall F1
@@
GPT-4o (Config 3, Epoch 1) $45$ $42$ $\mathbf{93.33\%}$ $80.77\%$ $\mathbf{0.87}$
@@ 4.5.4 Evaluation of Full Rule Correspondence
Experiment Model Generated Rules Matched Rules Precision Recall F1

### v2

@@ 2 Background
LLMs can be broadly divided into two categories: traditional LLMs and reasoning LLMs. Traditional LLMs, such as the GPT series by OpenAI, DeepSeek-V3 or similar models, are primarily optimized for fluent language generation and general-purpose tasks. Their strength lies in producing coherent and contextually appropriate text. However, their capabilities in logical reasoning and structured problem-solving are limited. Reasoning LLMs represent a newer generation of models that are explicitly designed to perform structured reasoning tasks more effectively. These models consider additional training objectives, architectural innovations, or fine-tuning procedures that enhance their ability to perform logical inference, complex decision-making, and consistent multi-step problem-solving. Reasoning LLMs aim not only at linguistic fluency but also at improved logical accuracy and reliability in formal contexts.
@@ 3 Methodology
    Controls randomness; higher values yield more diverse outputs [13]. Set to 0 to prioritize consistency over creativity [14].
@@ 3.1 Segmentation into Law Snippets
A key challenge in instructing the LLMs was determining the optimal length for law snippets: overly long snippets risked losing critical information during formalization, whereas overly short ones hindered atom reuse. To address this, we instructed the model to split enumerations containing more than two elements into separate law snippets while preserving shorter ones intact.
@@ 3.3 Evaluation
We evaluate the generated rules across six dimensions, each operationalized as a concrete question: completeness (Q1), syntactic (Q2) and semantic correctness (Q3), deontic modality accuracy (Q4), precondition appropriateness (Q5), and meaningfulness/reuse of atom names (Q6).
It is important to note that a single law snippet may lead to the generation of multiple rules. Furthermore, Q1 is assessed based on the law snippet as a whole, taking into account all rules derived from it. In contrast, Q2 through Q6 are evaluated individually for each generated rule. The questions are ordered such that earlier ones address more general and fundamental aspects of correctness, while later ones examine increasingly fine-grained details. Importantly, the evaluation follows a short-circuiting scheme: if for some rule $r$ a question Qi with $i\geq 2$ is evaluated as false, then all subsequent questions Qj with $j>i$ are not considered and are implicitly assigned the value false.
@@
Consider, for instance, the following formalization of law snippet 8.2.1.a.xiv:
@@
This initial check is crucial to prevent the LLM from achieving a high score merely by formalizing only the simplest aspects of a problem.
@@
Hence, the question would be answered with false, and no further checks performed. Note that this output stems from an early variation of the prompt. Such an error did not occur in later iterations.
@@
where $\text{Q1}(l)$ evaluates the overall satisfaction of Q1 across the law snippet $l$ (i.e., by considering all generated rules together). The success score $S(l)$ for an individual law snippet $l$ is then defined as:
@@ 4.1 Prompt Development
The prompt employed in our experiments was derived through a series of iterative refinements. Beginning with an initial two-shot learning prompt, successive modifications were made to enhance the clarity of the instructions and the quality of the generated outputs. These iterations involved the inclusion of additional guidance and examples to better align the model’s behavior with the desired output format. The final version utilizes a three-shot learning approach. Listing 1 provides the complete and final prompt used in our evaluation.
@@ 4.2.2 Providing Only Previously Formalized Atoms
In a second variant, previously extracted atom names were provided collectively in a single user message, rather than replicating the entire prior dialogue history. For each new law snippet, three messages were sent to the LLM: (1) a system prompt, (2) a user message listing previously formalized atom names (cf. Listing 2), and (3) a user message containing the new law snippet to be formalized.
@@
Listing 2: user message including previous atom names
@@ 4.3 Fine-Tuning
In the present study, fine-tuning experiments were conducted with GPT-4o. Although the proprietary nature of GPT-4o does not allow direct access to the model weights, OpenAI provides fine-tuning capabilities for non-reasoning models via its platform. Note that OpenAI’s reasoning models are not amendable to fine-tuning.
@@
Epochs 33 33 33
Batch Size 11 44 44
LR Multiplier 22 $1.5$ 11
@@
Fine-tuning resulted in an improved success score after a single epoch of training under Configurations 2 and 3, relative to the baseline performance of the non-fine-tuned GPT-4o. However, subsequent training epochs led to a decline in performance, indicative of overfitting.
@@ 4.4 Two-Stage Pipeline
In the first stage, atom names were extracted from the legal texts. To this end, the LLMs were instructed to identify atom names alongside brief textual descriptions (Listing 3). Three illustrative examples were provided within the prompt, thus applying a three-shot learning strategy.
@@ 8.1.1 A Supplier must take the following actions to
Listing 3: Prompt for atom extraction (2 examples omitted)
@@
In this section, we compare our experimental findings with the results presented by Dragoni et al. [8]. To ensure methodological consistency, we restrict the comparison to the lower branch evaluation reported in their study. For this purpose, we employ the standard metrics of Precision, Recall, and their harmonic mean, the F1-score, defined as follows:
$\begin{array}[]{lcl}\text{Precision}&=&\displaystyle\frac{\text{TP}}{\text{TP}+\text{FP}}=\displaystyle\frac{\text{Matched Items}}{\text{Generated Items}}\\ \text{Recall}&=&\displaystyle\frac{\text{TP}}{\text{TP}+\text{FN}}=\displaystyle\frac{\text{Matched Items}}{\text{Gold Standard Items}}\\ \text{F1}&=&2\cdot\displaystyle\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}\end{array}$
In this context, Items refers to either atoms or rules, depending on the evaluation. True positives (TP) are items generated by the LLM that are also found in the gold standard. False positives (FP) are generated items not present in the gold standard. False negatives (FN) are not explicitly observed. The denominator in the classical recall formula, $\text{TP}+\text{FN}$, serves as a proxy for the total number of positives in the labeled set. However, in our case, the gold standard itself is the labeled set and contains only positives. Therefore, the total number of positives is directly known and equals the number of items in the gold standard.
@@ 4.5.1 Evaluation of Term Identification
Experiment Model
Generated
Atoms
Matched
Atoms
Precision Recall F1
@@ 4.5.3 Identification of Rule Counterparts
Experiment Model
Generated
Rules
Matched
Rules
Precision Recall F1
@@
GPT-4o (Config 3, Epoch 1) $45$ $42$ $\mathbf{93.33\%}$ $7\% 0.87\mathbf{0.87}$
@@ 4.5.4 Evaluation of Full Rule Correspondence
Experiment Model
Generated
Rules
Matched
Rules
Precision Recall F1